# Gentle-CLIP: Exploring Aligned Semantic In Low-Quality Multimodal Data With Soft Alignment

### v1

@@ Abstract
Multimodal fusion breaks through the barriers between diverse modalities and has already yielded numerous impressive performances. However, in various specialized fields, it is struggling to obtain sufficient alignment data for the training process, which seriously limits the use of previously elegant models. Thus, semi-supervised learning attempts to achieve multimodal alignment with fewer matched pairs but traditional methods like pseudo-labeling are difficult to apply in domains with no label information. To address these problems, we transform semi-supervised multimodal alignment into a manifold matching problem and propose a new method based on CLIP, named Gentle-CLIP. Specifically, we design a novel semantic density distribution loss to explore implicit semantic alignment information from unpaired multimodal data by constraining the latent representation distribution with fine granularity, thus eliminating the need for numerous strictly matched pairs. Meanwhile, we introduce multi-kernel maximum mean discrepancy as well as self-supervised contrastive loss to pull separate modality distributions closer and enhance the stability of the representation distribution. In addition, the contrastive loss used in CLIP is employed on the supervised matched data to prevent negative optimization. Extensive experiments conducted on a range of tasks in various fields, including protein, remote sensing, and the general vision-language field, demonstrate the effectiveness of our proposed Gentle-CLIP.
1 Introduction
@@
As a pivotal foundation for numerous tasks[11, 10], multimodal learning has become the focus of many researches[8, 9]. By integrating information from diverse modalities such as texts, images and so forth, multimodal models can acquire more comprehensive information to enhance the generalization of the learned representations[36]. Meanwhile, multimodal fusion enables networks to emulate human-like multiple perceptual capabilities and address the inherent challenges like data scarcity, noise and ambiguity in various domains, from computer vision to healthcare[14, 3].
@@
In order to better harness the latent alignment information, most previous researches have primarily focused on the frameworks and the pretraining objectives, which can guide models to effectively understand the relationships between different modalities. Result from the sufficient development of image and text field, many studies have made great progress in the vision-language field[6, 51, 31, 33].Thereinto, CLIP employs a contrastive pretraining task [4] on large-scale datasets and gets robust multimodal representation. Due to simple framework and general pretraining task, it has good portability and scalability as well as competitive performance compared with supervised methods, hence it becomes the baseline for various vision-language works[53, 34]. Besides the aforesaid traditional field, in other intersecting domains, great breakthroughs have been also made by applying CLIP. EchoCLIP[13] improves the performance of cardiac imaging models by learning the association between cardiac ultrasound images and expert texts while ProtST[1] can capture more protein function information by aligning protein sequences and textual property descriptions. Moreover, in the field of zero-shot video recognition, Open-VCLIP[22] also shows excellent performance by leveraging the similar paradigm, which proves the powerful effects of CLIP.
@@
Figure 1: Comparison among CLIP, S-CLIP as well as Gentle-CLIP on how to adopt unpaired multimodal data. (a) CLIP only takes use of the matched data for multimodal fushion while ignoring the valuable information in unlabeled data. (b) S-CLIP attempts to improve the alignment performance by two novel pseudo-labeling methods but it limits to the language modality and heavily relies on the way how to measure similarity between samples. (c) Gentle-CLIP try to explore latent alignment from unmatched multimodal data based on the characteristics of data themselves, which needs less expert knowledge and can extend to various fields.
@@
Nonetheless, there are still many specialized fields where it is usually difficult to obtain sufficient alignment data[15] while traditional multimodal models like CLIP can only learn from matched pairs, which greatly limites the performance of previously elaborate models. In order to break the above dilemma, several studies paid attention to these specialized fields and attempted to reduce the number of matched pairs for pretraining[16, 17, 18, 19]. The main idea is to modify the loss function of CLIP and apply semi-supervised learning method[37, 38] to explore the latent alignment information in the unlabeled data. Recently, a new research improves the original CLIP and proposes S-CLIP[28] which introduces two novel pseudo-labeling methods for unlabeled images and achieve state-of-the-art in various specialized semi-supervised vision-language fields.
@@
However, pseudo-labeling learning methods may only be limited to the fields with class information and have difficulties scaling to other specialized multimodal domains where pseudo-labels are struggling to obtain. Meanwhile, the knowledge of generating pseudo-labels only relies on the insufficient labeled data, which leads to narrow ken and may loss much potential alignment information. In addition, the quality of pseudo-label has a great impact on the final performance so the learning process is unstable and even negative[39]. In order to solve these problems, it is necessary to design new semi-supervised methods for multimodalities, which can capture latent alignment information in unpaired data and be well extended to various multimodal domains.
@@
Therefore, we propose a novel semi-supervised learning method for multimodal alignment based on CLIP, named as Gentle-CLIP. We believe that ultimate representation is composed of modality, structure as well as semantic and the key to multimodal alignment is to capture the same semantic representation while ignoring the other two representations. On the premise of two modal data with the same semantic distribution, we design a new pretraining task based on manifold matching and design a novel loss called semantic density distribution(SDD) to better concentrate on the implicit alignment information among vast unpaired multimodal data. Moreover, we introduce multi-kernel maximum mean discrepancy(MK-MMD) to eliminate the difference between modality representations while self-supervised contrastive loss is used to prevent mode collapse and enhance the robustness of semantic distribution. At the same time, we apply contrastive loss from CLIP on the matched multimodal pairs to keep the correct learning direction. Gentle-CLIP tries to explore alignment relationship in latent space and it can be extended to various multimodal domains due to task irrelevance. Through end-to-end learning, the mutual constraints between losses prevent negative optimization and implicitly expand the knowledge range. Our approach attempts to achieve multimodal alignment from essence and it can be transferred to different multimodal frameworks[49, 50, 52]. The comparison between Gentle-CLIP and other learning strategies is shown in Figure 1.
@@
In short, our contributions are summarized as follows: (1) We contribute a groundbreaking perspective for the semi-supervised multimodal alignment problem by transforming it into a manifold matching problem, which brings a new pathway to exploit the implicit alignment information in the rich, yet largely unmatched multimodal data. (2) We design a novel semantic density distribution loss with less computing cost and it can be applied in various specialized fields as well as different multimodal frameworks. We introduce other objectives based on theoretical analysis about the components of representation and propose Gentle-CLIP to realize multimodal alignment with less supervised pairs. Moreover, our method can be applied to other domains with two-stream networks [35], such as knowledge distillation[47, 46], self-supervised learning[48] and domain adaptation. (3) We conduct extensive experiments in various fields and prove the effectiveness of our method. And Gentle-CLIP outperforms the existing semi-supervised methods in specialized fields because of larger knowledge range. We also explain the effects of key modules and provide a feasible usage paradigm for the specialized fields without enough supervised pairs.
2 Related Works
@@
Multimodal alignment. Multimodality enhances understanding and decision-making by integrating information from multiple sensorymodalities[81, 82]. Thereinto, ALBEF [77] aligns visual and language representations, using momentum distillation to improve multimodal embeddings. FLAVA [79] enhances multitask and cross-modal learning by jointly pretraining text and images. ALIGN [80] jointly trains language and image encoders, significantly enhancing performance across various vision and text benchmarks. In recent years, researches around CLIP has further optimized computational efficiency and model representation capabilities. For instance, FLIP[53] brings lower computation and faster training times by randomly removing a large number of image patches during training process while SoftCLIP[7] applies fine-grained interior self-similarity as a softening target for cross-modal learning to alleviate the strict mutual exclusion problem. Moreover, latent diffusion models[11] generates reliable text embeddings as condition by using pretrained text encoder of CLIP and CLIPSelf [12] enhances region-level representation through self-distillation from CLIP’s image encoder, which proves the powerful effects of CLIP.
@@
Semi-supervised learning. Semi-supervised learning [61, 69, 70] uses both labeled and unlabeled data to improve model training, encompassing strategies like pseudo-labeling [62, 73, 74], where models self-label their training data, and self-supervised learning [63, 64, 76, 75], which generates its own labels, thereby broadening its application in various fields. vONTSS [65] utilizes the von Mises-Fisher distribution and optimal transport for semi-supervised neural topic modeling to improve topic extraction in text datasets. SSGD [66] proposes a new semi-supervised domain generalization method that enhances model robustness under domain shifts through stochastic modeling and style augmentation. SS-ORL [68] employs a semi-supervised offline reinforcement learning approach, improving learning outcomes by utilizing unlabeled trajectories and limited complete action data. Semi-supervised learning ensures performance with fewer samples[71, 72], and the CLIP algorithm focuses on the complementarity and integration of different modalities. However, within the specialized domains, how to achieve these complementarity and integration across different modalities remains an issue that needs attention and resolution.
3 Method
3.1 Problem Description and Distribution Assumption
@@
Different from the general vision-lanuage field, there could be only limited available matched pairs between specific associated modalities while it is relatively simple to get a large amount of data with similar semantic distribution in each modality. Therefore, we propose Gentle-CLIP, which uses massive unmatched data as well as limited matched pairs to realize more generalized alignment through semi-supervised learning. Formally, for any two modalities $\bf A$ and $\bf B$, we employ a small number of matched pairs $\left\{a_{i},b_{i}\right\}_{i=1}^{N}$ and a large number of unmatched data $\left\{a_{j}\in\bf A\right\}_{j=1}^{M_{1}}$ as well as $\left\{b_{j}\in\bf B\right\}_{j=1}^{M_{2}}$ to train our model. Through sampling respectively from unpaired two sets, we acquire $\left\{a_{j}\in\bf A\right\}_{j=1}^{M}$ and $\left\{b_{j}\in\bf B\right\}_{j=1}^{M}$ as unsupervised training data which are only considered to have similar semantic distribution rather than strict one-to-one matching. Based on a natural assumption, models can be fully trained on adequate unmatched multimodal data for robust representation.
@@
Assumption 1 (Semantic Distribution Similarity Assumption, SDSA). We suppose that the latent embedding is a combination representation of modality, structure as well as semantic and more detailed analysis will be displayed in Appendix A. The goal of multimodal alignment is to find the same semantic representation and get rid of the interference from the other two representations. If the overall semantic distributions of $\left\{a_{j}\in\bf A\right\}_{j=1}^{M}$ and $\left\{b_{j}\in\bf B\right\}_{j=1}^{M}$ are similar, we can find a embedding space $\bf{S}\subseteq\bf{R}^{K}$ where $u_{j}$ and $v_{j}$ are the embedding representations respectively from $\bf{A}$ and $\bf{B}$. When the density distributions of $\left\{u_{j}\in\bf S\right\}_{j=1}^{M}$ as $\bf U$ and $\left\{v_{j}\in\bf S\right\}_{j=1}^{M}$ as $\bf V$ are similar, this space $\bf S$ is the semantic embedding space of $\bf A$ and $\bf B$. Consequently, when datasets from two modalities have the similar semantic distribution and their volumes are large enough, we can find the aligned semantic space by narrowing the gap between the density distribution from two modalities rather than strict matching relationship or pseudo-labeling method[45]. Through the above assumption, the semi-supervised multimodal alignment can be transformed into a manifold matching problem.
3.2 The Framework of Gentle-CLIP
@@
Figure 2 introduces the conceptual overview of Gentle-CLIP. Due to the convenience and efficiency of CLIP, our proposed method follows to design the two-stream network[35]. Each stream includes an encoder network $F_{i}(\cdot),i\in\{\bf A,\bf B\}$ and a projection head network $H_{i}(\cdot),i\in\{\bf A,\bf B\}$, which are applied to map the data from original space into embedding space. The network from different streams adopt different backbone with unshared weight. We introduce multi-kernel maximum mean discrepancy loss(MK-MMD) as well as self-supervised contrastive loss(SSL) and design a novel semantic density distribution loss(SDD) to learn potential alignment relationship in large amounts of unpaired data. Through contrastive metrix, we apply contrastive loss(CL) on limited supervised pairs to guarantee proper optimization of the model. Moreover, this framework could be extended to various modalities and the objectives in Gentle-CLIP can also migrate to other multimodal alignment works[31, 32]. A detailed description of loss functions will be shown below.
Figure 2: The overall framework of Gentle-CLIP. Thereinto, $N$ is the number of matched pairs while $M$ denotes the unlabeled scale. Based on CLIP, we redesign new objectives which work on latent space to acquire robust representation by exploring potential alignment from unmatched data.
3.3 Objective loss
Supervised Alignment Guidance: Based on problem description, there are limited matched pairs $\left\{a_{i},b_{i}\right\}_{i=1}^{N}$ and a large number of unmatched data $\left\{a_{j},b_{j}\right\}_{j=1}^{M}$. Due to the lack of sufficient data, we are supposed to learn generalized representations through unsupervised data and take advantage of explicit alignment relationship as ground-truth to achieve precise alignment. We apply contrastive loss $\mathcal{L}_{\mathrm{CL}}$ in CLIP, which is to maximize the representation similarity between matched pairs while minimize the similarity between negative pairs. The format of this loss function is shown as follows:
$\mathcal{L}_{\mathrm{CL}}=-\frac{1}{2n}\sum_{i=1}^{n}\left(\log\frac{\exp\left(\mathcal{S}(u_{i},v_{i})/\tau\right)}{\sum_{j=1}^{n}\exp\left(\mathcal{S}(u_{i},v_{j})/\tau\right)}+\log\frac{\exp\left(\mathcal{S}(v_{i},u_{i})/\tau\right)}{\sum_{j=1}^{n}\exp\left(\mathcal{S}(v_{i},u_{j})/\tau\right)}\right)$ (1)
where $n=\lfloor\frac{N}{M+N}B\rfloor$ and $B$ is the batch size. $u_{i}$ and $v_{i}$ are representations in latent space respectively from two modalities. $\tau$ is a learnable temperature parameter and $\mathcal{S}(\cdot,\cdot)$ denotes cosine similarity. We expect to apply supervised as well as unsupervised data in every batch to jointly train the model due to the reason that a mass of unsupervised data can bring richer alignment information while matched pairs could lead to more accurate learning.
@@ 3.3 Objective loss
Coarse-Grained Modality Adaptation: MK-MMD[44] is used to measure the gap between two probability distributions $P$ as well as $Q$ and the core idea of this method is that samples $\{p_{1},p_{2},\ldots,p_{m}\}$ as well as $\{q_{1},q_{2},\ldots,q_{n}\}$ drawn from $P$ and $Q$ should keep similar statistical properties if the two distributions are the same. Specifically, MK-MMD maps the data from original space to Reproducing Kernel Hilbert Spac(RKHS) by kernel functions[43] and we can compare the difference between distributions in this space. Through linear combination of multiple kernel functions, we could get a more robust mapping function to RKHS where we can easily distinguish two distributions even though they are similar in original space. The formula is shown as follows:
@@
Fine-grained Semantic Distribution Alignment: Since MK-MMD focuses on the whole distribution rather than sample level so it is imprecise and can only achieve macro alignment which is not enough for representation alignment. Consequently, we propose a new objective named as semantic density distribution loss(SDD) to explore more fine-grained information from unpaired data and realize more refined alignment. SDD is inspired from the perspective of probability density distribution estimation[29], hence it could keep an eye on specific sample representation while take the whole semantic distribution alignment into consideration at the same time. The formula is shown as follows:
@@
where $t_{i}$ is the sample from set $T$ and we apply sample variance with Bessel’s Correction. $\sigma(\cdot)$ can lead model to pay more attention to narrowing the gap between semantic distributions from different modalities while avoid close cluster. Through training, $\mathcal{L}_{\mathrm{SDD}}$ can make the semantic aligned data from different modalities keep the similar probability density distribution in latent space. Meanwhile, the time complexity of $\mathcal{L}_{\mathrm{SDD}}$ is $\mathcal{O}(B^{2})$ which is the same as $\mathcal{L}_{\mathrm{CL}}$. In addition, Sinkhorn algorithm can also achieve fine-grained alignment[84] but it brings huge computing costs and may not fit for low-quality multimodal data due to slashing alignment. More details will show in Appendix B.
Self-supervised Distribution Stability: Rely on self-supervised contrastive loss(SSL)[41, 40], we can adequately find out implicit information from single modality and get robust feature representation. In the field of multimodal alignment with limit matched pairs, we find that it is essential to apply this objective because it can pull away the representations of different samples in the latent space with incomplete alignment guidance. In other words, if SSL is not employed, the data without alignment constraint may gather into a tight cluster. To be specific, we apply augmentation to generate positive pairs and the format of $\mathcal{L}_{\mathrm{SSL}}$ is displayed as follows:
@@
where $z$ is latent embedding and $z_{i}^{+}$ denotes the representation of corresponding positive sample. In our method, each modality is supposed to apply this loss while $\mathcal{S}(\cdot,\cdot)$ is calculated by cosine similarity. We denote $u_{i}$ and $v_{i}$ as practical latent representation respectively from different two modalities and corresponding objectives are named as $\mathcal{L}_{\mathrm{SSL-U}}$ as well as $\mathcal{L}_{\mathrm{SSL-V}}$. According to the above constraint, we propose new loss named as $\mathcal{L}_{GC}$ and the formula is shown as follows:

### v2

# Set-CLIP: Exploring Aligned Semantic From Low-Alignment Multimodal Data Through A Distribution View

@@ Abstract
Multimodal fusion breaks through the boundaries between diverse modalities and has already achieved notable performances. However, in many specialized fields, it is struggling to obtain sufficient alignment data for training, which seriously limits the use of previously effective models. Therefore, semi-supervised learning approaches are attempted to facilitate multimodal alignment by learning from low-alignment data with fewer matched pairs, but traditional techniques like pseudo-labeling may run into troubles in the label-deficient scenarios. To tackle these challenges, we reframe semi-supervised multimodal alignment as a manifold matching issue and propose a new methodology based on CLIP, termed Set-CLIP. Specifically, by designing a novel semantic density distribution loss, we constrain the latent representation distribution with fine granularity and extract implicit semantic alignment from unpaired multimodal data, thereby reducing the reliance on numerous strictly matched pairs. Furthermore, we apply coarse-grained modality adaptation and unimodal self-supervised guidance to narrow the gaps between modality spaces and improve the stability of representation distributions. Extensive experiments conducted on a range of tasks in various fields, including protein analysis, remote sensing, and the general vision-language field, validate the efficacy of our proposed Set-CLIP method. Especially with no paired data for supervised training, Set-CLIP is still outstanding, which brings an improvement of $144.83\%$ over CLIP.
Introduction
@@
As a pivotal foundation for numerous tasks(Rombach et al. 2022; Zhou et al. 2023), multimodal learning has become the focus of many research(Rasekh et al. 2024; Bhalla et al. 2024). By integrating information from diverse modalities such as texts, images and more, multimodal models can derive more comprehensive information to enhance the generalization of the learned representations(Gao et al. 2020). Meanwhile, such fusion enables networks to emulate human-like multiple perceptual capabilities and address the inherent challenges like data scarcity, noise and ambiguity in various domains, from computer vision to healthcare(Wang et al. 2022; Zang et al. 2024).
@@
To better harness latent alignment information, previous studies have mainly concentrated on developing frameworks and pretraining objectives to enhance multimodal understanding. Result from the sufficient developments of images and texts, many studies have made great progress in the vision-language field(Radford et al. 2021; Chen et al. 2022; Kim, Son, and Kim 2021).Thereinto, CLIP employs a contrastive pretraining task on large-scale datasets and gets robust multimodal representation. Due to effective framework and general pretraining task, it has good portability and competitive performance compared with supervised methods, hence it becomes the baseline for various vision-language works(Li et al. 2023). Besides the aforesaid traditional field, in other intersecting domains, great breakthroughs have been also made by applying CLIP. EchoCLIP(Christensen et al. 2024) improves the performance of cardiac imaging models by correlating ultrasound images with expert texts while ProtST(Xu et al. 2023a) capture more protein function information by aligning protein sequences and textual property descriptions. Moreover, in the field of zero-shot video recognition, Open-VCLIP(Weng et al. 2023) also shows excellent performance by leveraging the similar paradigm, which proves the powerful effects of CLIP.
@@
Figure 1: Comparison among CLIP, S-CLIP and Set-CLIP on how to adopt unpaired multimodal data. (a) CLIP only uses the matched data for multimodal fusion while ignores the valuable information in unlabeled data. (b) S-CLIP attempts to improve the alignment performance by two pseudo-labeling losses but it limits itself to the language modality and heavily relies on the way how to measure similarity between samples. (c) Set-CLIP tries to explore more latent information from unmatched multimodal data by fine-grained distribution alignment, which is based on data themselves without much expert knowledge.
@@
Nonetheless, there are still many specialized fields where it is usually difficult to obtain sufficient alignment data(Li et al. 2022a) while traditional multimodal models like CLIP can only learn from matched pairs, which greatly limites the performance of previously elaborate models. In order to break above dilemma, several studies paid attention to these specialized fields and attempted to learn from low-alignment data with fewer matched pairs for pretraining(Zhou et al. 2022; Mu et al. 2022). The main idea is to modify the loss function of CLIP and apply semi-supervised learning method(Yang et al. 2022) to explore the latent alignment information from unlabeled data. Recently, a new research improves the original CLIP and proposes S-CLIP(Mo et al. 2023) which introduces two novel pseudo-labeling losses for unlabeled images and achieve state-of-the-art in various specialized semi-supervised vision-language fields.
@@
However, pseudo-labeling methods may only be limited to the fields with class information and have difficulties scaling to other specialized multimodal domains where pseudo-labels are struggling to obtain. Meanwhile, the knowledge of generating pseudo-labels only relies on the insufficient labeled data, which leads to narrow ken and may loss much potential alignment information. In addition, the quality of pseudo-label has a great impact on the final performance so the learning process is unstable and even negative(Arazo et al. 2020). In order to solve these problems, it is necessary to design new semi-supervised methods for multimodalities, which can capture latent alignment information in unpaired data and be well extended to various multimodal domains.
@@
Therefore, we propose a novel semi-supervised learning method for multimodal alignment based on CLIP, named as Set-CLIP. We believe that ultimate representation is composed of modality, structure as well as semantic and the key to multimodal alignment is to capture the same semantic representation while ignoring the other two parts. On the premise of two aligned modal data with the same semantic distribution, we design a new pretraining task based on manifold matching and a novel loss called semantic density distribution(SDD) to better concentrate on the implicit alignment among vast unpaired multimodal data. Moreover, we introduce multi-kernel maximum mean discrepancy(MK-MMD) to eliminate the gap between modality representations while self-supervised contrastive loss is used to prevent mode collapse and enhance the robustness of semantic representation. At the same time, we apply contrastive loss from CLIP on the matched multimodal pairs to keep the correct learning direction. Set-CLIP tries to explore alignment relationship in latent space and it can be extended to various multimodal domains due to task irrelevance. Through end-to-end learning, the mutual constraints between losses prevent negative optimization and implicitly expand the knowledge range. Our approach can be transferred to different multimodal frameworks(Li et al. 2022b, c) and the comparison of Set-CLIP with other strategies is shown in Figure 1.
@@
In short, our contributions are summarized as follows: (1) We contribute a groundbreaking perspective for the semi-supervised multimodal alignment problem by reframing it as a manifold matching problem, which brings a new pathway to exploit the implicit alignment information in the rich, yet largely unmatched multimodal data. (2) We design a novel semantic density distribution loss with fine-grained constrain and it can be applied in various specialized fields as well as different multimodal frameworks. We introduce other objectives based on theoretical analysis about the components of representation and propose Set-CLIP to realize multimodal alignment with less supervised pairs. Moreover, our method can be applied to other domains with two-stream networks(Cao, Lu, and Zhang 2024), such as knowledge distillation(Pham et al. 2022), self-supervised learning(He et al. 2020) and domain adaptation. (3) We conduct extensive experiments in various fields and prove the advantages of Set-CLIP. Moreover, We also explain the effects of key modules and provide a feasible usage paradigm for the specialized fields with limited supervised pairs.
Related Works
@@
Multimodal alignment. Multimodality enhances understanding and decision-making by integrating information from multiple sensorymodalities(Martin et al. 2022). Thereinto, ALBEF (Li et al. 2021) aligns visual and language representations, using momentum distillation to improve multimodal embeddings. FLAVA (Singh et al. 2022) enhances multitask and cross-modal learning by jointly pretraining text and images. ALIGN (Jia et al. 2021) jointly trains language and image encoders, significantly enhancing performance across various vision and text benchmarks. In recent years, researches around CLIP has further optimized computational efficiency and model representation capabilities. For instance, FLIP(Li et al. 2023) brings lower computation and faster training times by randomly removing a large number of image patches during training process while SoftCLIP(Gao et al. 2024) applies fine-grained interior self-similarity as a softening target to alleviate the strict mutual exclusion problem. Moreover, latent diffusion models generates reliable text embeddings as condition by using pretrained text encoder of CLIP and CLIPSelf (Wu et al. 2023) enhances region-level representation through self-distillation from CLIP’s image encoder, which proves the powerful effects of CLIP.
@@
Semi-supervised learning. Semi-supervised learning (Van Engelen and Hoos 2020) uses both labeled and unlabeled data to improve training process, encompassing strategies like pseudo-labeling (Cascante-Bonilla et al. 2021), where models self-label their training data, and self-supervised learning (Krishnan, Rajpurkar, and Topol 2022; Liu et al. 2022), which explores the values in data itself. vONTSS (Xu et al. 2023b) utilizes the von Mises-Fisher distribution and optimal transport for semi-supervised neural topic modeling to improve topic extraction in text datasets. SSGD (Zhou, Loy, and Liu 2023) proposes a new semi-supervised domain generalization method that enhances model robustness under domain shifts through stochastic modeling and style augmentation. SS-ORL (Zheng et al. 2023b) employs a semi-supervised offline reinforcement learning approach, improving learning outcomes by utilizing unlabeled trajectories and limited complete action data. Semi-supervised learning ensures performance with fewer samples, but in the specialized domains, how to achieve complementarity and integration across different modalities with limited paired data remains an issue that needs attention and resolution(Zong, Mac Aodha, and Hospedales 2024).
Method
Problem Description and Assumption
@@
Different from the general vision-language field, there could be only limited available matched pairs between specific associated modalities while it is relatively simple to get a large amount of unimodal data with similar semantic distribution. Therefore, we propose Set-CLIP, which uses massive unmatched data as well as limited matched pairs to realize more generalized alignment through semi-supervised learning. Formally, for any two modalities $\bf A$ and $\bf B$, we employ a small number of matched pairs $\left\{a_{i},b_{i}\right\}_{i=1}^{N}$ and a large number of unmatched data $\left\{a_{j}\in\bf A\right\}_{j=1}^{M_{1}}$ as well as $\left\{b_{j}\in\bf B\right\}_{j=1}^{M_{2}}$ to train our model. Through sampling respectively from two unpaired sets, we acquire $\left\{a_{j}\in\bf A\right\}_{j=1}^{M}$ and $\left\{b_{j}\in\bf B\right\}_{j=1}^{M}$ as unsupervised training data which are only considered to have similar semantic distribution rather than strict one-to-one matching. Based on a natural assumption, models can be trained on these adequate unmatched multimodal data.
@@
Assumption 1 (Semantic Distribution Similarity Assumption, SDSA). We suppose that the latent embedding is a combination representation of modality, structure as well as semantic and more detailed analysis will be displayed in Appendix A. The goal of multimodal alignment is to find the same semantic representation and get rid of the interference from the other two representations. If the overall semantic distributions of $\left\{a_{j}\in\bf A\right\}_{j=1}^{M}$ and $\left\{b_{j}\in\bf B\right\}_{j=1}^{M}$ are similar, we can find a embedding space $\bf{S}\subseteq\bf{R}^{K}$ where $u_{j}$ and $v_{j}$ are the embedding representations respectively from $\bf{A}$ and $\bf{B}$. When the density distributions of $\left\{u_{j}\in\bf S\right\}_{j=1}^{M}$ as $\bf U$ and $\left\{v_{j}\in\bf S\right\}_{j=1}^{M}$ as $\bf V$ are similar, this space $\bf S$ is the semantic embedding space of $\bf A$ and $\bf B$. Consequently, when datasets from two modalities have the similar semantic distribution and their volumes are large enough, we can find the aligned semantic space by narrowing the gap between the density distribution from two modalities rather than strict matching relationship or pseudo-labeling method. Through the above assumption, we can explore the value from unpaired data and the semi-supervised multimodal alignment can be transformed into a manifold matching problem.
The Framework of Set-CLIP
@@
Figure 2 introduces the conceptual overview of Set-CLIP. Due to the convenience and efficiency of CLIP, our proposed method follows to design the two-stream network. Each stream includes an encoder network $F_{i}(\cdot),i\in\{\bf A,\bf B\}$ and a projection head network $H_{i}(\cdot),i\in\{\bf A,\bf B\}$, which are applied to map the data from original space into embedding space. The network from different streams adopt different backbone and is trained from scratch. We introduce MK-MMD as well as self-supervised contrastive loss(SSL) and design a novel semantic density distribution loss(SDD) to learn potential alignment in large amounts of unpaired data. Through contrastive metrix, we apply contrastive loss(CL) on limited supervised pairs to guarantee proper optimization. The multimodal batch with the size of $B$ is composed of $\lfloor\frac{N}{M+N}B\rfloor$ paired data from $\left\{a_{i},b_{i}\right\}_{i=1}^{N}$ while the rest data is sampling from two unsupervised training datasets. A detailed description of loss functions will be shown below.
Figure 2: The overall framework of Set-CLIP. Here, $N$ is the number of matched pairs, while $M$ denotes the unlabeled scale. Pink indicates the loss objectives. Thereinto, MK-MMD is used to narrow the gap between the distribution spaces of different modalities, and SDD operates to reduce the differences between latent distributions in a fine-grained way. Applying SSL can enhance the robustness of representation, and the contrastive loss in CLIP ensures the proper optimization direction.
Objective Loss
@@
Coarse-Grained Modality Adaptation: There may be large discrepancy between the embedding distributions from different modalities. While the latent representations are supposed to have the similar distribution space when they become aligned, which can be treated as a domain adaptation problem. Thereinto, MK-MMD(Long et al. 2015) is used to measure the gap between two probability distributions $P$ and $Q$ while the core idea of this method is that samples $\{p_{1},p_{2},\ldots,p_{m}\}$ as well as $\{q_{1},q_{2},\ldots,q_{n}\}$ drawn from $P$ and $Q$ should keep similar statistical properties if the two distributions are the same. Specifically, MK-MMD maps the data from original space to Reproducing Kernel Hilbert Spac(RKHS) by kernel functions(Roth and Steinhage 1999) and we can compare the difference between distributions in this space. Through linear combination of multiple kernel functions, we could get a more robust mapping function to RKHS where we can easily distinguish two distributions even though they are similar in original space. The formula is shown as follows:
@@
Fine-grained Semantic Distribution Alignment: Since MK-MMD pays attaention to the whole distribution rather than sample level so it is imprecise and can only achieve macro alignment which is not enough for representation alignment. Consequently, we propose a novel objective named as semantic density distribution loss(SDD) to explore more fine-grained information from unpaired data and realize more refined alignment. SDD is inspired from the perspective of probability density distribution estimation, hence it could keep an eye on specific sample representation while take the whole semantic distribution alignment into consideration at the same time. The formula is shown as follows:
@@
where $t_{i}$ is the sample from set $T$ and we apply sample variance with Bessel’s Correction. $\sigma(\cdot)$ can lead model to focus on narrowing the gap between semantic distributions while avoid close cluster. By employing $\mathcal{L}_{\mathrm{SDD}}$, semantic aligned data from different modalities will get similar density distribution in the latent space during training. Meanwhile, the time complexity of $\mathcal{L}_{\mathrm{SDD}}$ is $\mathcal{O}(B^{2})$ which is the same as $\mathcal{L}_{\mathrm{CL}}$. More details of SDD will show in Appendix B.
Supervised Alignment Guidance: Based on problem description, there are limited matched pairs $\left\{a_{i},b_{i}\right\}_{i=1}^{N}$ and a large number of unmatched data $\left\{a_{j},b_{j}\right\}_{j=1}^{M}$. Due to the lack of sufficient data, we are supposed to learn generalized representations through unsupervised data and take advantage of explicit alignment relationship as ground-truth to achieve precise alignment. We apply contrastive loss $\mathcal{L}_{\mathrm{CL}}$ in CLIP, which is to maximize the representation similarity between matched pairs while minimize the similarity between negative pairs. The format of this loss is shown as follows:
$\begin{split}\mathcal{L}_{\mathrm{CL}}=-\frac{1}{2n}\sum_{i=1}^{n}(&\log\frac{\exp\left(\mathcal{S}(u_{i},v_{i})/\tau\right)}{\sum_{j=1}^{n}\exp\left(\mathcal{S}(u_{i},v_{j})/\tau\right)}\\ +&\log\frac{\exp\left(\mathcal{S}(v_{i},u_{i})/\tau\right)}{\sum_{j=1}^{n}\exp\left(\mathcal{S}(v_{i},u_{j})/\tau\right)})\end{split}$ (6)
where $n$ denotes the paired size in a batch and is generally $\lfloor\frac{N}{M+N}B\rfloor$. $u_{i}$ and $v_{i}$ are representations in latent space respectively from two modalities. $\tau$ is a learnable temperature parameter and $\mathcal{S}(\cdot,\cdot)$ denotes cosine similarity. We expect to apply supervised as well as unsupervised data in every batch to jointly train the model due to the reason that a mass of unsupervised data can bring richer alignment information while matched pairs could lead to more accurate learning.
Self-supervised Distribution Stability: Rely on self-supervised contrastive loss(SSL)(Chen et al. 2020; Gao, Yao, and Chen 2021), we can adequately find out implicit information from single modality and get robust feature representation. In the field of multimodal alignment with limit matched pairs, we find that it is essential to apply this objective because it can pull away the representations of different samples in the latent space with incomplete alignment guidance. In other words, if SSL is not employed, the data without alignment constraint may gather into a tight cluster. To be specific, we apply augmentation to generate positive pairs and the format of $\mathcal{L}_{\mathrm{SSL}}$ is displayed as follows:
@@
where $z$ is latent embedding and $z_{i}^{+}$ denotes the representation of corresponding positive sample. In our method, each modality is supposed to apply this loss while $\mathcal{S}(\cdot,\cdot)$ is calculated through cosine similarity. We denote $u_{i}$ and $v_{i}$ as practical latent representation respectively from different two modalities and the corresponding objectives are named as $\mathcal{L}_{\mathrm{SSL-U}}$ as well as $\mathcal{L}_{\mathrm{SSL-V}}$. According to the above constraint, we propose new loss named as $\mathcal{L}_{GC}$ and the formula is shown as follows: